What wilk be the electric potentialif the source charge is +4.02x10^-13C and a test charge will be placed 4.55x10^-12m from it?


1
Expert's answer
2021-03-21T11:19:53-0400

"\phi=k\frac{q}{r}=9\cdot10^9\cdot\frac{4.02\cdot10^{-13}}{4.55\cdot10^{-12}}\approx795\ (MV)" . Answer